# SkySense: A Multi-Modal Remote Sensing Foundation Model Towards Universal Interpretation for Earth Observation Imagery

### 15 Dec 2023 v1

@@ 1 Introduction
Remote Sensing Imagery (RSI) interpretation is crucial in understanding our common home, the Earth [17, 70], via quite diverse tasks [82, 14, 48, 5], e.g. crop monitoring, natural disaster management, etc. Every task may require significant dedicated efforts and resources to build a task-specific model. Recently, Foundation Model emerges as a pre-trained generic model that excels in a wide range of downstream tasks [86]. Hence, there is a soaring interest in exploring a comprehensive Remote Sensing Foundation Model (RSFM) for many Earth Observation (EO) tasks.
The key question naturally arises: What is essential for a RSFM? First of all, an ideal RSFM should possess the ability to perceive multi-modal temporal RSI. EO heavily relies on multi-modal time series remote sensing data, including temporal optical and Synthetic Aperture Radar (SAR) data. Individual modality offers unique advantages and complements to each other. For example, optical images provide rich spectral bands and texture details but are susceptible to weather [87]. In contrast, SAR sensors capture clear imagery in all weather conditions [35, 42]. Moreover, the time series of such data provide the crucial temporal clue to various tasks [25, 5, 83] like change prediction. Second, a RSFM should be easy to tailor when being deployed for EO tasks using different modalities (i.e., single- and multi-modal) at different spatial (i.e., pixel-, object-, and image-level) granularities. Last but not the least, remote sensing data is inherently contingent on their space-time coordinates, which provide rich regional and seasonal geo-context that benefits RSI interpretation a lot, as indicated in [43, 12, 28, 36]. Therefore, a RSFM shall bear the vital capability of effective geo-context learning and utilization.
@@
Previous works on RSFM [2, 51, 1, 76, 19, 64, 72, 8, 66, 75, 53, 50, 4, 56, 52, 73, 37] have demonstrated their preliminary success on several specific datasets. However, most of these studies [2, 51, 1, 76, 64, 72, 8, 66, 53, 50, 56, 52, 73] primarily employ optical RGB imagery for model pre-training, which limits their effectiveness in tasks that require multispectrum and more modalities [61]. Methods like Scale-MAE, CROMA, etc. [56, 53, 64, 52, 24, 74] are built on static RSI, missing the critical temporal clue for the tasks like crop mapping. In addition, models such as SatLas and RVSA [4, 72, 24, 74] do not integrate the geo-context that is vital for various RSI tasks [43, 12, 28, 36]. These RSFMs, while proficient in certain areas, are limited in their application to EO tasks, as they primarily address only specific aspects of RSI interpretation.
In this paper, we propose SkySense, a billion-scale Multi-Modal Remote Sensing Foundation Model (MM-RSFM). SkySense incorporates 2.06 billion parameters and is pre-trained on a large-scale multi-modal dataset which comprises 21.5 million RSI temporal sequences extracted from high-spatial-resolution optical images (HSROIs), medium-resolution temporal multispectral imagery (TMsI) and temporal SAR imagery (TSARI). To handle the multi-modal temporal RSI sequences, SkySense employs a factorized multi-modal spatiotemporal encoder to perform spatial feature extraction and multi-modal temporal fusion independently, since RSI sequence is spatially-aligned in nature. It leads to a modular design allowing flexible use of its modules, i.e., the spatial encoder can be either used alone or in combination of fusion module to support tasks from static single-modal to temporal multi-modal. This design delivers strong modeling of RSI sequences while using substantially less parameters compared to common 3D structures [85, 49]. The factorized encoder is pre-trained by Multi-Granularity Contrastive Learning to construct features from different modal and spatial granularities. Furthermore, we propose Geo-Context Prototype Learning to generate regional prototypes from RSI features given geo-locations. This approach enhances multi-modal spatiotemporal representation learning by leveraging the regional context clue hidden in numerous unlabeled RSI.
@@
  • We extensively compare SkySense with 18 recently published RSFMs. Our model has achieved the SOTA performance, supparssing the latest models like GFM, SatLas and Scale-MAE by over $2.5\%$ on average. We hope the release of pre-trained weights will contribute to the Remote Sensing community and facilitate future works.
@@ 2 Related Work
2.1 Vision Foundation Model
The research focus of Vision Foundation Model has evolved from supervised learning on extensive labeled data [29, 71, 44, 21] to the recent Contrastive Learning paradigm [30, 11, 27, 55, 7] for unsupervised pre-training on vast amounts of unlabeled data. The unprecedented success of Large Language Models [22, 69] has inspired Masked Image Modeling (MIM) methods (e.g. MAE [31], BEiT [3], etc.) emerging as another paradigm.
@@ 2.2 Remote Sensing Foundation Model
Remote sensing data inherently integrates space-time coordinates and has diverse spatial scales. The maintstream RSFMs extend recent foundation model techniques to space-time RS data, such as Contrastive Learning. For instance, GASSL [2] utilized geo-location prediction as an additional pre-text task in the MoCo-v2 framework [13]. Multiple views with different sizes were utilized by DINO-MC [76] for self-supervised learning within the DINO framework [7]. SeCo [51] and CACo [50] both proposed Contrastive Learning to perceive short-term and long-term changes by using the spatiotemporal structure of temporal RSI sequences. Besides, there are works either improving the MIM-based framework [64, 72, 56] or exploring the model scale-up [8]. For example, RingMo [64] modified MAE to adapt to the dense objects in RSI. SatMAE [19] employed TMsI to enhance the performance on temporal sequences. Scale-MAE [56] built a framework with scale-aware encoder. Recent efforts such as CMID [53] and GFM [52] have commenced to explore amalgamation of CL and MIM strategies. Concurrently, CROMA [24] and DeCUR [74] investigated multi-modal pre-training for single- and multi-modal tasks using static imagery. In this study, we propose a comprehensive MM-RSFM, SkySense, to fill the gap between existing RSFMs, i.e., single modality of RingMo, CACo, etc., static input of Scale-MAE, CROMA, etc., and geo-context neglection of SatLas, RVSA, etc.
@@ 3.1 Pre-training Dataset
We curate an extensive multi-modal remote sensing dataset with temporal sequences, containing RSI from various sources: HSROIs from WorldView-3, 4, etc. (RGB band), TMsI from Sentinel-2 (B2-8, B8A, B11-12 band) and TSARI from Sentinel-1 (VV, VH Polarization). All data is geo-spatially aligned. Strictly speaking, HSROIs and TMsI shall be categorized to the optical modality, while TSARI falls to the SAR modality. However, due to HSROIs and TMsI’s significant difference in spectral band and ground sample distance, we regard HSROIs and TMsI as two distinct modalities for simplicity in this paper. The dataset comprises 21.5 million training samples, each consisting of a static HSROI with rich texture details, a TMsI containing temporal and multispectral data, a TSARI providing backscatter polarization under cloud coverage, and the metadata like geo-location and acquisition date for geo-context modeling. This dataset covers a great variety of scenarios across resolution, spectrum, and imaging mechanism. More details of the data are included in the supplementary materials. We construct the input for SkySense as $\left\{x_{HR},x_{Ms},x_{SAR}\right\}$, where $x_{HR}$ represents a static HSROI; $x_{Ms}$ is a Sentinel-2 TMsI after filtering cloudy images, where we randomly select 20 images to form the sequence; and $x_{SAR}$ stands for a standard-calibrated TSARI, from whom we randomly select 10 images for training.
@@ 3.2 Model Architecture
Factorized Multi-Modal Spatiotemporal Encoder. The overall architecture of our method is illustrated in Fig. 2. In a multi-modal input $\left\{x_{HR},x_{Ms},x_{SAR}\right\}$, pixels within each RSI naturally align with the others given the same geo-location. Upon this, we propose a factorized encoder that initially extracts spatial features from each RSI independently and then fuses them to capture a multi-modal spatiotemporal representation. The design separates spatial feature extraction from the feature fusion, enabling the integration of the clues from modality, time and geo-context.
@@
Attentional Geo-Context Integration. Each RSI’s geographical location is linked to rich region-specific geo-context. It is valuable for RSI interpretation as indicated by [43, 12, 28, 36]. To utilize this contextual clue to enhance $F_{\mathbf{fus}}^{\text{mm}}$, we employ a region-specific prototype set $\mathcal{P}\in\mathbb{R}^{N_{R}\times N_{p}\times d}$ (shown on the right side of Fig. 2), where $N_{R}$ is the number of regions, $N_{p}$ represents the number of prototypes for each region and $d$ denotes the feature dimension. The learning procedure of $\mathcal{P}$ will be elaborated in Sec. 3.3. Specifically, a regional prototype subset $\mathcal{P}_{r}\in\mathbb{R}^{N_{p}\times d}$ is chosen from $\mathcal{P}$ based on the geo-location embedded with $F_{\mathbf{fus}}^{\text{mm}}$. $F_{\mathbf{fus}}^{\text{mm}}$ is then attended to the prototypes of $\mathcal{P}_{r}$ through the attention mechanism, as shown in Eq. 9. The weights, computed from $\operatorname{Softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)$, facilitate a soft selection of prototypes in accordance with their similarity to $F_{\mathbf{fus}}^{\text{mm}}$. The final representation $F_{\mathbf{fus}}\in\mathbb{R}^{N_{S}\times 2d}$ is generated by concatenation of $F_{\mathbf{fus}}^{\text{mm}}$ and weighted sum of prototypes from $\mathcal{P}_{r}$ along the feature dimension. The prototypes represent a set of discriminative features linked to certain semantics like water body, cropland, etc. By finding the similar ones to $F_{\mathbf{fus}}^{\text{mm}}$, we provide the standard representations of certain semantics to complement $F_{\mathbf{fus}}^{\text{mm}}$,
@@ 3.3 Pre-training
$F_{i}^{\mathbf{obj}}\in\mathbb{R}^{N_{C}\times d}$ denotes as object-level feature generated from unsupervised clustering on pixel-level feature vectors $f_{i}^{\mathbf{pix}}$ in a single RSI, where $N_{C}$ is the number of clusters. The clustering employs the same Sinkhorn-Knopp algorithm [6] we apply for Geo-Context Prototype Learning, as shown later. $f_{i}^{\mathbf{obj}}\in\mathbb{R}^{d}$ is the vector representing cluster center in $F_{i}^{\mathbf{obj}}$, which can be viewed as a general representation for a set of collected $f_{i}^{\mathbf{pix}}$. It normally links to a certain ground object or semantics. The object-level contrastive learning loss is computed as Eq. 12,
@@
Unsupervised Geo-Context Prototype Learning. Different regions characterize distinct geographic landscapes and seasonal dynamics [36, 34] due to disparities in topography and climate. Prior arts have shown that the enlarged context can benefit the RSI interpretations [43, 12, 28, 36]. In this work, $F_{\mathbf{fus}}^{\text{mm}}$ captures rich spatiotemporal clue for a small area. By clustering on numerous $F_{\mathbf{fus}}^{\text{mm}}$, higher-level regional semantics are obtained as implicit geo-knowledge for a vast geo-spatial scope (see Fig. 5). Thus, we propose Geo-Context Prototype Learning to unsupervisedly extract regional geo-context from $F_{\mathbf{fus}}^{\text{mm}}$ during pre-training.
@@ 4.2 Performance on Single-Modal Tasks
We evaluate SkySense on 4 representative single-modal tasks. All experiments are conducted using consistent fine-tuning settings for fairness. The supplementary materials include implementation details and visualization results.
@@ 5 Discussions & Ablation Studies
Multi-modal Pre-training Effectiveness. In addition to confirming the effectiveness of using multi-modal data in downstream tasks, we investigate the impact of multi-modal pre-training on single-modal tasks, compared with pre-training on fewer modalities. We conduct experiments on iSAID for static HSROI segmentation and fMoW-S2 for temporal multispectral classification. Two versions of the pre-trained model are tested: a simple version pre-trained only with optical imagery (HSROIs, TMsI), and SkySense, which includes HSROIs, TMsI, and TSARI for pre-training. The rest settings remain consistent. The results in Tab. 4(a) show that SkySense consistently outperforms the simple version, suggesting that the introduction of SAR data benefits representation learning of other modalities. This may attribute to the implicit clue brought by SAR data through Cross-Modal Alignment. It provides another perspective on the necessity of SkySense’s multi-modal pre-training.
@@
What does Geo-Context Prototype (GCP) Learn? We utilize the Dyna.-MM dataset for experiment as it contains diverse geo-locations worldwide. For segmentation task in Tab. 4a, adding GCP in downstream task leads to a further gain of 0.5% mIoU compared to the strong multi-modal baseline (v). Moreover, a comparison between (i) and (ii) shows a 0.5% mIoU improvement using GCP for single-modal task. It demonstrates GCP’s consistent performance gain in single- and multi-modal scenarios.
@@ 6 Conclusion & Future Work
In this paper, we present SkySense, a large-scale MM-RSFM for interpretation of EO imagery. SkySense allows using its modules flexibly to accommodate different scenarios and consistently outperforms other models on a variety of tasks, showcasing its exceptional generalization ability and strong performance. We hope SkySense will inspire further research on MM-RSFM and its release may contribute to sustainable innovations thriving in the Remote Sensing community. Our future work aims to explore the model’s scalability, extend its applications to more EO tasks and improve the compatibility for other satellite sources.
Acknowledgement
This work was supported by Ant Group Research Fund.

### 22 Mar 2024 v2

@@ 1 Introduction
Remote Sensing Imagery (RSI) interpretation is crucial in understanding our common home, the Earth [17, 70], via quite diverse tasks [84, 14, 49, 5], e.g. crop monitoring, natural disaster management, etc. Every task may require significant dedicated efforts and resources to build a task-specific model. Recently, Foundation Model emerges as a pre-trained generic model that excels in a wide range of downstream tasks [88, 82]. Hence, there is a soaring interest in exploring a comprehensive Remote Sensing Foundation Model (RSFM) for many Earth Observation (EO) tasks.
The key question naturally arises: What is essential for a RSFM? First of all, an ideal RSFM should possess the ability to perceive multi-modal temporal RSI. EO heavily relies on multi-modal time series of remote sensing data, including temporal optical and Synthetic Aperture Radar (SAR) data. Individual modality offers unique advantages and complements to each other. For example, optical images provide rich spectral bands and texture details but are susceptible to weather [89]. In contrast, SAR sensors capture clear imagery in all weather conditions [34, 42]. Moreover, the time series of such data provide the crucial temporal clue to various tasks [24, 5, 85] like change prediction. Second, a RSFM should be easy to tailor when being deployed for EO tasks using different modalities (i.e., single- and multi-modal) at different spatial (i.e., pixel-, object-, and image-level) granularities. Last but not the least, remote sensing data is inherently contingent on their space-time coordinates, which provide rich regional and seasonal geo-context that benefits RSI interpretation a lot, as indicated in [44, 12, 27, 35, 43]. Therefore, a RSFM shall bear the vital capability of effective geo-context learning and utilization.
@@
Previous works on RSFM [2, 52, 1, 77, 19, 64, 72, 8, 66, 75, 54, 51, 4, 57, 53, 73, 37] have demonstrated their preliminary success on several specific datasets. However, these RSFMs, while proficient in certain areas, are limited in their applications to EO tasks, due to factors such as single-modal pre-training and the neglect of geo-context.
In this paper, we propose SkySense, a billion-scale Multi-Modal Remote Sensing Foundation Model (MM-RSFM). SkySense incorporates 2.06 billion parameters and is pre-trained on a large-scale multi-modal dataset which comprises 21.5 million RSI temporal sequences extracted from high-spatial-resolution optical images (HSROIs), medium-resolution temporal multispectral imagery (TMsI) and temporal SAR imagery (TSARI). To handle the multi-modal temporal RSI sequences, SkySense employs a factorized multi-modal spatiotemporal encoder to perform spatial feature extraction and multi-modal temporal fusion independently, since RSI sequence are spatially-aligned in nature. It leads to a modular design allowing flexible use of its modules, i.e., the spatial encoder can be either used alone or in combination of the fusion module to support tasks from static single-modal to temporal multi-modal. This design delivers strong modeling of RSI sequences while using substantially less parameters compared to common 3D structures [87, 50]. The factorized encoder is pre-trained by Multi-Granularity Contrastive Learning to construct features from different modal and spatial granularities. Furthermore, we propose Geo-Context Prototype Learning to generate regional prototypes from RSI features given geo-locations. This approach enhances multi-modal spatiotemporal representation learning by leveraging the regional context clue hidden in numerous unlabeled RSI.
@@
  • We extensively compare SkySense with 18 recently published RSFMs. Our model has achieved the SOTA performance, supparssing the latest models like GFM, SatLas and Scale-MAE by over $2.5\%$ on average. We hope the release of pre-trained weights will contribute to the Remote Sensing community and facilitate future research.
@@ 2.1 Remote Sensing Foundation Model
Recent Remote Sensing Foundation Models draw their primary inspiration from the research on Vision Foundation Model [28, 71, 45, 21, 29, 11, 26, 56, 7, 30, 3]. Remote sensing data inherently integrates space-time coordinates and has diverse spatial scales. The maintstream RSFMs extend the foundation model techniques to space-time RS data, such as Contrastive Learning. For instance, GASSL [2] utilized geo-location prediction as an additional pre-text task in the MoCo-v2 framework [13]. Multiple views with different sizes were utilized by DINO-MC [77] for self-supervised learning within the DINO framework [7]. SeCo [52] and CACo [51] both proposed Contrastive Learning to perceive short-term and long-term changes by using the spatiotemporal structure of temporal RSI sequences. Besides, there are works either improving the MIM-based framework [64, 72, 57] or exploring the model scale-up [8]. For example, RingMo [64] modified MAE to adapt to the dense objects in RSI. SatMAE [19] employed TMsI to enhance the performance on temporal sequences. Scale-MAE [57] built a framework with scale-aware encoder. Recent efforts such as CMID [54] and GFM [53] have commenced to explore amalgamation of CL and MIM strategies. Concurrently, CROMA [23] and DeCUR [74] investigated multi-modal pre-training for single- and multi-modal tasks using static imagery. In this study, we propose a comprehensive MM-RSFM, SkySense, to fill the gap in existing RSFMs, i.e., single modality of RingMo, CACo, etc., static input of Scale-MAE, CROMA, etc., and the neglect of geo-context of SatLas, RVSA, etc.
@@ 3.1 Pre-training Dataset
We curate an extensive multi-modal remote sensing dataset with temporal sequences, containing RSI from various sources: HSROIs from WorldView-3, 4, etc. (RGB band), TMsI from Sentinel-2 (B2-8, B8A, B11-12 band) and TSARI from Sentinel-1 (VV, VH Polarization). All data is geo-spatially aligned. Strictly speaking, HSROIs and TMsI shall be categorized to the optical modality, while TSARI falls to the SAR modality. However, due to HSROIs and TMsI’s significant difference in spectral band and ground sample distance, we regard HSROIs and TMsI as two distinct modalities for simplicity in this paper. The dataset comprises 21.5 million training samples, each consisting of a static HSROI with rich texture details, a TMsI containing temporal and multispectral data, a TSARI providing backscatter polarization under cloud coverage, and the metadata like geo-location and acquisition date for geo-context modeling. This dataset covers a great variety of scenarios across resolution, spectrum, and imaging mechanism. More details of the data are included in the supplementary materials. We construct the input for SkySense as $\left\{x_{HR},x_{Ms},x_{SAR}\right\}$, where $x_{HR}$ represents a static HSROI; $x_{Ms}$ is a Sentinel-2 TMsI after filtering cloudy images, where we randomly select 20 images to form the sequence; and $x_{SAR}$ stands for a standard-calibrated TSARI, from which we randomly select 10 images for training.
@@ 3.2 Model Architecture
Factorized Multi-Modal Spatiotemporal Encoder. The overall architecture of our method is illustrated in Fig. 2. In a multi-modal input $\left\{x_{HR},x_{Ms},x_{SAR}\right\}$, the pixels within each RSI naturally align with the others given the same geo-location. Upon this, we propose a factorized encoder that initially extracts spatial features from each RSI independently and then fuses them to capture a multi-modal spatiotemporal representation. The design separates spatial feature extraction from the feature fusion, enabling the integration of the clues from modality, time and geo-context.
@@
Attentional Geo-Context Integration. Each RSI’s geographical location may reveal rich region-specific geo-context. It is valuable for RSI interpretation as indicated by [44, 12, 27, 35]. To utilize this contextual clue to enhance $F_{\mathbf{fus}}^{\text{mm}}$, we employ a region-specific prototype set $\mathcal{P}\in\mathbb{R}^{N_{R}\times N_{p}\times d}$ (shown on the right side of Fig. 2), where $N_{R}$ is the number of regions, $N_{p}$ represents the number of prototypes for each region and $d$ denotes the feature dimension. The learning procedure of $\mathcal{P}$ will be elaborated in Sec. 3.3. Specifically, a regional prototype subset $\mathcal{P}_{r}\in\mathbb{R}^{N_{p}\times d}$ is chosen from $\mathcal{P}$ based on the geo-location embedded with $F_{\mathbf{fus}}^{\text{mm}}$. $F_{\mathbf{fus}}^{\text{mm}}$ is then attended to the prototypes of $\mathcal{P}_{r}$ through the attention mechanism, as shown in Eq. 9. The weights, computed from $\operatorname{Softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)$, facilitate a soft selection of prototypes in accordance with their similarity to $F_{\mathbf{fus}}^{\text{mm}}$. The final representation $F_{\mathbf{fus}}\in\mathbb{R}^{N_{S}\times 2d}$ is generated by concatenation of $F_{\mathbf{fus}}^{\text{mm}}$ and weighted sum of prototypes from $\mathcal{P}_{r}$ along the feature dimension. The prototypes represent a set of discriminative features linked to certain semantics like water body, cropland, etc. By finding the similar ones to $F_{\mathbf{fus}}^{\text{mm}}$, we provide the standard representations of certain semantics to complement $F_{\mathbf{fus}}^{\text{mm}}$,
@@ 3.3 Pre-training
$F_{i}^{\mathbf{obj}}\in\mathbb{R}^{N_{C}\times d}$ denotes object-level feature generated from unsupervised clustering on pixel-level feature vectors $f_{i}^{\mathbf{pix}}$ in a single RSI, where $N_{C}$ is the number of clusters. The clustering employs the same Sinkhorn-Knopp algorithm [6] we apply for Geo-Context Prototype Learning, as shown later. $f_{i}^{\mathbf{obj}}\in\mathbb{R}^{d}$ is the vector representing the cluster centers in $F_{i}^{\mathbf{obj}}$, which can be viewed as a general representation for a set of collected $f_{i}^{\mathbf{pix}}$. It usually corresponds to a certain ground object or semantics. The object-level contrastive learning loss is computed as Eq. 12,
@@
Unsupervised Geo-Context Prototype Learning. Different regions characterize distinct geographic landscapes and seasonal dynamics [35, 33] due to disparities in topography and climate. Prior arts have shown that the enlarged context can benefit the RSI interpretations [44, 12, 27, 35]. In this work, $F_{\mathbf{fus}}^{\text{mm}}$ captures rich spatiotemporal clues for a small area. By clustering on numerous $F_{\mathbf{fus}}^{\text{mm}}$, higher-level regional semantics are obtained as implicit geo-knowledge for a vast geo-spatial scope (see Fig. 5). Thus, we propose Geo-Context Prototype Learning to unsupervisedly extract regional geo-context from $F_{\mathbf{fus}}^{\text{mm}}$ during pre-training.
@@ 4.2 Performance on Single-Modal Tasks
We evaluate SkySense on 4 representative single-modal tasks. All experiments are conducted using consistent fine-tuning settings for fairness. The supplementary materials include implementation details, visualization results and additional experiments on frozen backbone tuning.
@@ 5 Discussions & Ablation Studies
Multi-modal Pre-training Effectiveness. In addition to confirming the effectiveness of using multi-modal data in downstream tasks, we investigate the impact of multi-modal pre-training on single-modal tasks, compared with pre-training on fewer modalities. We conduct experiments on iSAID for static HSROI segmentation and fMoW-S2 for temporal multispectral classification. Two versions of the pre-trained model are tested: a simple version pre-trained only with optical imagery (HSROIs, TMsI), and SkySense, which includes HSROIs, TMsI, and TSARI for pre-training. The rest of the settings remain consistent. The results in Tab. 4(a) show that SkySense consistently outperforms the simple version, suggesting that the introduction of SAR data benefits representation learning of other modalities. This may attribute to the implicit clue brought by SAR data through Cross-Modal Alignment. It provides another perspective on the necessity of SkySense’s multi-modal pre-training.
@@
What does Geo-Context Prototype (GCP) Learn? We utilize the Dyna.-MM dataset for experiment as it contains diverse geo-locations worldwide. For the segmentation task in Tab. 4a, adding GCP in downstream tasks leads to a further gain of 0.5% mIoU compared to the strong multi-modal baseline (v). Moreover, a comparison between (i) and (ii) shows a 0.5% mIoU improvement using GCP for the single-modal task. It demonstrates GCP’s consistent performance gain in single- and multi-modal scenarios.
@@ 6 Conclusion & Future Work
In this paper, we present SkySense, a large-scale MM-RSFM for interpretation of EO imagery. SkySense allows using its modules flexibly to accommodate different scenarios and consistently outperforms other models on a variety of tasks, showcasing its exceptional generalization ability and strong performance. We hope SkySense will inspire further research on MM-RSFM and its release may contribute to sustainable innovations thriving in the Remote Sensing community. As part of our future work, we plan to incorporate the language modality, thereby extending SkySense’s applications to more EO tasks.
7 Acknowledgment
This work was in part supported by the National Natural Science Foundation of China under Grants 42030102 and 42371321.